# Neural Network Reprogrammability: A Unified Theme on Model Reprogramming, Prompt Tuning, and Prompt Instruction

### v1

@@ 4 Emergent Insights from the Reprogrammability Perspective
Our unified framework (Sec. 2) and taxonomy (Sec. 3) do more than just catalogue existing methods. By showing how this perspective illuminates some interesting phenomena, e.g., in-context learning and chain-of-thought reasoning, as well as the trade-offs inherent in different adaptation choices, we can gain insights into how these techniques relate to fundamental properties of the pre-trained model itself.
@@ 4.2 Implications of Taxonomic Position
Interface $\text{Interface Sensitivity}(f,\lambda)\approx\begin{cases}\epsilon&\text{if }|f|<S_{\text{threshold}}\\ \beta\cdot\log(|f|/S_{\text{threshold}})&\text{otherwise }\end{cases}$
where $|M|$ represents model size, $S_{\text{threshold}}$ is the emergence threshold, and $\beta$ is a scaling factor. This behavior dictates distinctly different adaptation strategies for models above and below certain scale thresholds, as a sublinear relationship between model size and optimal manipulation volume is observed (Lester et al., 2021). Specifically, the proportion of parameters needed for effective adaptation decreases as models grow larger, following an empirically observed relationship, such that $\text{Optimal Prompt Size}\propto\text{Model Size}^{\alpha}$, where $\alpha<1$. Such a sublinear scaling (Lester et al., 2021) suggests that larger models develop more parameter-efficient representational capabilities.
In addition, as model scale increases, the complexity of required output alignment operations decreases (Chowdhery et al., 2023). Larger pre-trained models can directly generate properly formatted outputs with minimal post-processing, effectively internalizing output alignment functions that would require explicit implementation in smaller models.
@@ 5.1 Towards a Theory of Reprogrammability
  • Role of Pre-training Objectives: How do different self-supervised or supervised pre-training objectives (e.g., masked language modeling (Devlin et al., 2019), contrastive learning (Radford et al., 2021), next-token prediction (Radford et al., 2019)) influence a model’s reprogrammability for diverse downstream tasks?
@@
Answering these questions could profoundly inform how we design and train future foundation models to be inherently more adaptable, efficient, and broadly reprogrammable.

### v2

@@ 4 Emergent Insights from the Reprogrammability Perspective
Our framework (Sec. 2) and taxonomy (Sec. 3) do more than just catalogue existing methods. By showing how this perspective illuminates some interesting phenomena, e.g., in-context learning and chain-of-thought reasoning, as well as the trade-offs inherent in different adaptation choices, we can gain insights into how these techniques relate to fundamental properties of the pre-trained model itself.
@@ 4.2 Implications of Taxonomic Position
Interface $\text{Interface Sensitivity}(f,\lambda)\approx\begin{cases}\epsilon&\text{if }|f|<S_{\text{threshold}}\\ \beta\cdot\log(|f|/S_{\text{threshold}})&\text{otherwise, }\end{cases}$
where $|M|$ represents model size, $S_{\text{threshold}}$ is the emergence threshold, and $\beta$ is a scaling factor. This behavior dictates distinctly different adaptation strategies for models above and below certain scale thresholds, as a sublinear relationship between model size and optimal manipulation volume is observed (Lester et al., 2021). Specifically, the proportion of parameters needed for effective adaptation decreases as models grow larger, following an empirically observed relationship, such that $\text{Optimal Prompt Size}\propto\text{Model Size}^{\alpha}$ with $\alpha<1$. This sublinear scaling (Lester et al., 2021) suggests that larger models develop more parameter-efficient representational capabilities.
In addition, as model scale increases, the complexity of required output alignment operations decreases (Chowdhery et al., 2023). For large-scale deployment across related tasks, deeper manipulations with sample-specific components offer the strongest performance, whilst for targeted adaptation to a specific task with a smaller model, fully fine-tuning may remain competitive despite its parameter inefficiency.
@@ 5.1 Towards a Theory of Reprogrammability
  • Role of Pre-training Objectives: How do different self-supervised or supervised pre-training objectives (e.g., masked language modeling (Devlin et al., 2019), contrastive learning (Radford et al., 2021), next-token prediction (Radford et al., 2019)) influence the pre-trained model’s reprogrammability for diverse downstream tasks?
@@
Answering these questions could profoundly inform how we design and train future foundation models to be inherently more efficient and broadly reprogrammable.